I have tried to figure this out all night. Even some of my peers tried and none of us could do it. Can you please help me?
0.2(s+0.2)+0.5(s-0.4)=5.44 the answer is 8 solve for s

Found 3 solutions by Heretohelpya!, queenofit, askmemath:
Answer by Heretohelpya!(2)
Answer by queenofit(120)

First you must remember that fractions and decimals seem to be one of the most difficult things for most people, so don't feel discouraged.
Let's work your problem together OK?
You gave the equation as
0.2(s+0.2)+0.5(s-0.4)=5.44 you know the answer should be s=8
the first thing we do is distribute the numbers in front of the ( ) throughout the ( ) by multiplication. Use your calculator along with me and see how I got the following from your equation.
your first section was 0.2(s+0.2) this means you mult. the 0.2 in front of the ( to all the numbers and variables inside the (). in this case it would be
0.2*s and 0.2*0.2 that becomes 0.2s+0.04
the second part is 0.5(s-0.4) multiply this out and you get 0.5s-0.2
now your new equation after we do all the multiplication is
0.2s+0.04+0.5s-0.2=5.44
do you see how I got this?
now you must put the common factors together.
0.2s+0.5s=0.7s and 0.04-0.02=-.16
so your new equation is
0.7s-0.16=5.44
but you must put the numbers on one side and the variables on the other. the new equation will be
0.7s=5.44+0.16 you know to change the signs, I am sure. then add
0.7s=5.6 now you must divide by 0.7 to solve for s
s=8
I hope you can follow me, if you need further assistance, feel free to ask....
hope this helps.
A.

Answer by askmemath(368)

0.2(s+0.2)+0.5(s-0.4)=5.44
In order to remove the brackets,we multiply each term inside the brackets with the term outside the bracket.
0.2s + 0.04 + o.5s - 0.2 = 5.44
0.7s -0.16 = 5.44
Now keep all the terms containing s on one side and take the other terms to the right side. Remember to change the sign as you do.
0.7s = 5.44 + 0.16
0.7s = 5.6
Dividing by 0.7 throughout
s = 8